OpenStudy (anonymous):

In a right triangle, the long leg is three times the length of the short leg. The hypothenuse is the square of the short leg. Find the exact perimeter and area of this triangle.

OpenStudy (anonymous):



OpenStudy (anonymous):

P=x+3x+x^2=x^2+4 A=1/2*x*3x=1/2*3x^2

OpenStudy (anonymous):

P=x^2+4x, typo

OpenStudy (anonymous):

but the directions say to find the EXACT perimeter and area...so there shouldn't be any variable...correct? Don't we have to use something like SOH-CAH-TOA?

OpenStudy (anonymous):

unless you know how long one of the legs is, you can't solve the area and perimeter without variables

OpenStudy (anonymous):

you can find the angles, but that won't help you with the area/perimeter

OpenStudy (anonymous):

hmm...idk regardless, thanks for your time :)

OpenStudy (anonymous):

yw